OpenStudy (anonymous):

Prove the identity, cos3(theta)= 4cos^3(theta)-3cos(theta)

OpenStudy (anonymous):

Use deMoivre Theorem! It's piece of cake!

OpenStudy (anonymous):

lo,l maybe for you but looking at that theorem just confused me even more..

OpenStudy (anonymous):

DO you know binomial theorem?

OpenStudy (anonymous):

no, not yet.

OpenStudy (anonymous):

\[\cos3x+i*\sin3x=\cos^3x +3*\cos^2x*i*\sin x+3*\cos x * i^2 \sin^2 x +i^3 * \sin^3 x\]\[\cos3x=\cos^3 x - 3* \cos x * \sin^2 x\]\[\cos3x= \cos^3x - 3 \cos x(1-\cos^2 x)\]\[\cos3x=4\cos^3x - 3 \cos x\]

OpenStudy (anonymous):

Understand? Now give me a best response and close the question! ;D

OpenStudy (anonymous):

lmao, you got it!!

OpenStudy (anonymous):

thanks !!